OpenStudy (anonymous):

am i solving this correctly? limx->0 sin 7x/4x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \sin7x }{ 4x }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ 4 }{ 7 }( \frac{ \sin x }{ x })\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 7 }\lim_{x \rightarrow 0} \frac{ \sin x }{ x }\]

OpenStudy (anonymous):

sin x/x is just 1 and 4/7 times 1 is just 4/7

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\sin ax}{ax}=1\] So, \[\lim_{x\to0}\frac{\sin7x}{4x}=\frac{7}{7}\lim_{x\to0}\frac{\sin7x}{4x}=\frac{7}{4}\lim_{x\to0}\frac{\sin 7x}{7x}=\cdots \]

OpenStudy (anonymous):

The answer should be the reciprocal of what you got: 7/4.

OpenStudy (anonymous):

haha yes, ofcourse, typo there. i meants 7/4

OpenStudy (anonymous):

meant*

OpenStudy (anonymous):

so

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ 7 }{ 4 }( \frac{ \sin x }{ x })\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }\lim_{x \rightarrow 0} \frac{ \sin x }{ x }\]

OpenStudy (anonymous):

= 7/4

OpenStudy (anonymous):

No, multiply by 7/7 (=1) and then switch out the 4x for 7x, then the sin7x or 7x turns into one. 1 times 7/4 will be 7/4. You got the right answer, but in the wrong way.

OpenStudy (anonymous):

you just switch.. ok then